OpenStudy (anonymous):

Contagious diseases can spread very quickly. Suppose five people are ill during the first week of an epidemic and that each person who is ill spreads the disease to four people by the end of the next week. By the end of the tenth week of the epidemic, how many people have been affected by the illness?

OpenStudy (anonymous):

Well, let's take it one week at a time. How many will be sick at the end of week 1?

OpenStudy (anonymous):

ok theres 5 people ill in each spreads to 4 so 4*5=20 plus the 5=25

OpenStudy (anonymous):

Good =) Keep goin' Week 2

OpenStudy (anonymous):

25 * 10=250 people in 10 weeks right

OpenStudy (dumbcow):

think of it as exponential function, with time in terms of weeks you need initial value when t=0 you need to know the scale factor...after each week the amount multiplies by a constant each time

OpenStudy (anonymous):

Nono, it doesn't go like that.

OpenStudy (anonymous):

Dumbcow, I'm getting there =) Pattern recognition.

OpenStudy (anonymous):

Week 0: 5 people Week 1: 25 people Week 2?

OpenStudy (dumbcow):

sorry..i like to jump ahead to formula, then fill in the missing pieces

OpenStudy (anonymous):

I don't like to just hand a student a formula without making it mean something to them first.

OpenStudy (anonymous):

ok so 25 *4

OpenStudy (anonymous):

Brenda, you correctly said that after a week there would be 25 sick people. Now, in a weeks time, each of those 25 people will spread the disease to 4 new people.

OpenStudy (anonymous):

Good =) 25*4, but then we keep that same 25, so it's 25*4 + 25 which is just 25*5

OpenStudy (anonymous):

And I'd also point out that the first time, you did 5*4 + 5 which is just 5*5

OpenStudy (anonymous):

ok so 25*4=100+25=125

OpenStudy (anonymous):

so 125*4=500+125=625

OpenStudy (anonymous):

Mhmm Week 0: 5 Week 1: 25 (5*5) Week 2: 125 (5*5*5) Week 3: 625 (5*5*5*5)

OpenStudy (anonymous):

625*4=2500+625=3125 3125*4=12500+3125=15625 right

OpenStudy (anonymous):

You keep multiplying the number by 4 and then adding the number. I'd rather you just multiply by 5.

OpenStudy (anonymous):

Do you see that it's the same? n*4 + n = n*5

OpenStudy (anonymous):

15625*4=62500+15625=78125 78125*4=312500+78125=390625

OpenStudy (anonymous):

390625*4=1562500+390625=1953125

OpenStudy (anonymous):

are u sure this is right numbers are too big

OpenStudy (anonymous):

Okay, Brenda, what you're doing is right, and it's going to get you there. But there's an easier way, especially with the problems that do more than 10 weeks.

OpenStudy (anonymous):

And I'm sure it's right. The point is that it is very big numbers. It's called exponential growth.

OpenStudy (anonymous):

But let's slow down and see if we can notice the pattern. First, can you agree that instead of multiplying the number by 4 and then adding the number, we can just multiply by 5? Like for week 1, we start with 5 people. You did 5*4 + 5 = 25 Couldn't you just have done 5*5 = 25

OpenStudy (anonymous):

so what will be the answer 48828125

OpenStudy (anonymous):

If I tell you that's the correct answer, are you going to leave and not let me teach you the easy way to do this kind of problem?

OpenStudy (anonymous):

no im not

OpenStudy (anonymous):

Your answer is correct =)

OpenStudy (anonymous):

bye im leaving lol im just playing

OpenStudy (anonymous):

Okay, now here's the key to this kind of problem. Notice that there's a pattern to how the number of people is growing. What is the pattern? Week 0: 5 Week 1: 25 Week 2: 125 Week 3: 625 What is the pattern? How do I get from one week to the next?

OpenStudy (anonymous):

It might not be that obvious to you, but it multiplies by 5 every time.

OpenStudy (anonymous):

*5

OpenStudy (anonymous):

it mark my answer wrong are u sure thats it

OpenStudy (anonymous):

Right. And there's a whole family of functions that work exactly like that. To get to the next number, you multiply by something each time. It might double every time. It might triple every time. It might multiply by 10 every time. Or it could be getting smaller It could halve every time. It could be a third of the size each time. It might be one tenth every time.

OpenStudy (anonymous):

I think it only marked you wrong because of a misunderstanding of how the problem was written. I think you did one extra week.

OpenStudy (anonymous):

can u check

OpenStudy (anonymous):

Take your answer and divide by 5. That's what you should have put in, I believe.

OpenStudy (anonymous):

9765625

OpenStudy (anonymous):

Does it accept that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay, so the wording of this problem is kind of... lame. And I think it confused us mostly about how many times this disease was multiplying. Here's the important info: five people are ill during the first week each person who is ill spreads the disease to four people by the end of the next week Week 1: 5 Week 2: 25 Week 3: 125 Week 4: 625 Week 5: 5^5 ... Week 10: 5^10 = 9765625

OpenStudy (anonymous):

it marks it wrong

OpenStudy (anonymous):

Lol, well then it's somehow a mis-communication of the problem.